# thanks: equal contributionthanks: equal contribution

### v1

@@ Abstract
Nature uses elongated shapes and filaments to build stable structures, generate motion, and allow complex geometric interactions. In this Review, we examine the role of biological filaments across different length scales. From the molecular scale, where cytoskeletal filaments provides a robust but dynamic cellular scaffolding, over the scale of cellular appendages like cilia and flagella, to the scale of filamentous microorganisms like cyanobacteria which are among the most successful genera on Earth, and even to the scale of elongated animals like worms and snakes, whose motility modes inspire robotic analogues. We highlight the general mechanisms that couple form and function. Physical principles, such as classical elasticity and the non-reciprocity of active matter can be used to trace unifying themes linking these systems spanning about six orders of magnitude in length.
@@ I Introduction
The interconnectedness between form and function in biology has been quantified at least since D’Arcy Wentworth Thompson’s influential work Thompson (1917). Because geometrical features influence and direct biological activity through physical forces, it seems worthwhile to explore the link between geometry and biological function. In this Review, we examine a simple geometrical shape –the filament – which is pervasive in the architecture of living organisms, and discuss the function of filamentous shapes and the physical advantages they offer.
@@
Filamentous shapes permeate every length scale in the biological world. Natural evolution has found that the geometrical properties of long aspect-ratio shapes can be used in a multiplicity of useful ways. The biopolymers forming the cytoskeleton provide structural integrity to the cell and form an intracellular transportation system. Cytoskeletal networks generate cytokinesis Eggert et al. (2006), migration Paluch et al. (2006); Lämmermann et al. (2008), and apical contraction Martin et al. (2009), and even generate cytoplasmic streaming in the Drosophila oocyte Dutta et al. (2024).
@@
At length scales of a few microns, cilia and flagella can generate fluid flow and active motion due to their fluid dynamical interactions Lighthill (1976); Gilpin et al. (2020); Lauga and Powers (2009) and they can synchronize to generate metachronal waves Vilfan and Jülicher (2006); Gilpin et al. (2020). Microorganisms such as filamentous fungi form complex networks of hyphae following available water and resources, and can move cytoplasm along the network Klein and Paschke (2004). Filamentous fungi are also the group of fungi most often used in biotechnological processes and include perhaps the most important fungus in the history of humanity: Penicillium rubens, which was used by Alexander Fleming to extract penicillin Fleming (1929, 1941); Houbraken et al. (2011). The slime mold Physarum polycephalum performs cytoplasmic flow across its network akin to a generalised form of peristalsis, which is optimised when the wavelength of the peristaltic wave is of the order of the size of the network Alim et al. (2013); this filamentous network can even encode memory about nutrient location in its morphology Kramar and Alim (2021). In the prokaryotic domain, cyanobacteria were among the first microorganisms to straddle the transition from unicellular to multicellular organisms Grosberg and Strathmann (2007); Claessen et al. (2014), and include numerous filamentous species.
Box I | Geometry of filaments A filament can be described as a portion of a three-dimensional curve, whose local orientation changes smoothly along its length. The curve needs to be sufficiently smooth, that is, it is at least thrice continuously differentiable. The Frenet frame or trihedron is a local, right-handed, orthogonal coordinate system with basis $(\bm{t},\bm{n},\bm{b})$. We indicate the curve by $\bm{r}(s)$, where $\bm{r}$ is the position vector and $s$ is the arc length of the curve. The tangent unit vector to the curve is defined as $\bm{t}\equiv\frac{d\bm{r}}{ds}$. The normal unit vector $\bm{n}\equiv\frac{d\bm{t}}{ds}/|\frac{d\bm{t}}{ds}|$, and always satisfies $\bm{n}\cdot\bm{t}=0$. Furthermore, $\kappa\equiv|\frac{d\bm{t}}{ds}|$ is the local curvature of the curve $\bm{r}(s)$. Finally, the binormal vector $\bm{b}\equiv\bm{t}\times\bm{n}$ together with $\bm{t}$ and $\bm{n}$ form an orthonormal basis for space. These three vectors obey the Frenet–Serret equations $\frac{d\bm{t}}{ds}=\kappa\bm{n}\,,\quad\frac{d\bm{n}}{ds}=-\kappa\bm{t}+\tau\bm{b}\,,\quad\frac{d\bm{b}}{ds}=-\tau\bm{n}\,,$ where $\tau$ is the torsion. The Frenet trihedron can be useful in the description of the helical motion of free-swimming organisms Crenshaw and Edelstein-Keshet (1993); Friedrich and Jülicher (2009); Bente et al. (2020), sperm cells Jikeli et al. (2015); Friedrich and Jülicher (2007), chiral active particles Sevilla (2016), and for gliding motility Lettermann et al. (2024). [Uncaptioned image]
@@
Figure 1: Filamentous shapes in biology. (a) Growing complexity of filaments across orders of magnitude in length, from nucleic acids that store information, to biopolymers that provide structural integrity to cells and form networks spanning the entire cytoskeleton. Flagella and cilia self-assemble from polymers to generate motion. Bacterial cells such as cyanobacteria self-organize into filamentous shapes. Animals, from worms and nematodes to snakes and eels utilize elongated shapes to achieve locomotion and collective behaviour. (b) Hypothetical schematic diagram of the biological function emerging from the three main properties of biological filaments: their length, their rigidity (measured by, e.g., Young’s modulus), and their structural complexity, such as twisting, cross-linking into a network, up until the tissue level of the muscle fibers or an entire body.
@@
At much larger length scales, organisms such as worms, snakes, and eels have elongated bodies that grant them ecological advantages. Worms, for example, can form entangled ‘blobs’ that exhibit shear-thinning rheology Deblais et al. (2020) and reversible self-assembly Patil et al. (2023). Such advantages can be combined to create higher levels of complexity, such as in a network built from cross-linking filaments, as illustrated in Fig. 1.
@@
Although the systems described above span about nine orders of magnitude in size (from nucleic acids in the nanometre range in diameter, to snakes and eels in the metre range), it is possible to identify and discuss general physical features of elongated shapes (from macromolecules to whole organisms) that depend on the geometrical form of filaments and, in turn, influence their biological activity.
@@
How do cytoskeletal filaments generate structural integrity in the cell, how do groups of cilia generate fluid flow, or how do large elongated organisms generate undulatory locomotion? Understanding how physical systems interact via simple rules to generate interesting collective behaviour can and will have novel applications. The condensed-matter motto of ‘More is Different’ proposed by Philip Anderson Anderson (1972) is a guiding principle in this field. For example, a self-organizing robot exhibits behaviour intermediate between solid and liquid with adaptive capabilities such that it can flow over obstacles Saintyves et al. (2024). This combination of flexibility and structural integrity is reminiscent of a cell’s cytoskeleton, or even of snake locomotion. Each of these filamentous systems can serve as inspiration for new biosynthetic materials Fladung et al. (2024). Examples of this principle are the realisation of strain-stiffening metamaterials using PDMS slats that mimic cell behaviour Taale et al. (2023), or the use of carbon nanotubes to study the effect of crowding and filament stiffness on their microscopic dynamics Fakhri et al. (2010, 2009).
Physically, the rigidity and length of different biofilaments determine the class of models used to describe their mechanical and dynamical properties. If fluctuations are not important (because of the filament’s rigidity or because of the length scale of observation), the biofilaments can be effectively described as smooth lines. In this context, we can mathematically describe filaments as a curve in a three-dimensional (3D) space whose local orientation changes smoothly along their length. Box I describes the Frenet–Serret equations used in differential geometry to characterize a curve.
@@
At smaller length scales, such as inside the cell, the polymeric nature of biofilaments is consequential and this character is amplified by the nonequilibrium nature of the fluctuations. For example, variations in polymer stiffness and connectivity of actomyosin networks by actin filament bundling and cross-linking proteins can systematically tune the mechanical response of actin networks from extensile to contractile Stam et al. (2017); Banerjee et al. (2020).
Another dimension, useful in classifying filaments in the natural world, is the complexity of their organisation. Single globular actin proteins (G-actins) assemble in filaments (F-actin), which together with cross-linking proteins and a myriad of other molecules form cytoskeletal networks. Multiple cells, in turn, can assemble into simple filaments, such as the Oscillatoria lutea cyanobacteria, or form even more complex filamentous organisms, such as nematodes, eels and snakes.
@@
In this Review, we discuss the physical processes and properties of filamentous shapes in the biological world across length scales. Section II discusses the role of biofilaments inside the cell. Section  III reviews the role of filamentous cellular appendages, such as the locomotor function of cilia and flagella. In Sec. IV we discuss the biological properties of filamentous bacteria. Section V deals with macroscopic elongated organisms, such as snakes and eels. Finally, in Sec. VI we outline a roadmap of future work and outstanding questions at this exciting frontier shared by biology and physics, to establish a common framework describing the physics of biological filaments (Fig. 1)
@@ II Biopolymers & cytoskeletal filaments
The cytoskeleton is a 3D polymer-based scaffold that spans the entire interior of a cell and defines cell morphology Huber et al. (2013). The basic building blocks are few key proteins that assemble together into a broad range of filaments Fletcher and Mullins (2010). These filaments together with a host of proteins that bind to the filaments’ sides or ends form a network that provides mechanical stability to the cell Pegoraro et al. (2017). Using only a few building blocks allows great flexibility and economy of materials. For example, cytoskeletal polymers achieve the same rigidity (elastic modulus) of gel-like substances with two orders of magnitude less material Kollmannsberger and Fabry (2011); Pegoraro et al. (2017).
Biological filaments are subject to different types of deformations: extensional and compressional forces and moments that cause the filaments to bend and twist. Any deformation mode costs energy, which is continuously used in the cytoskeleton to maintain the cell in a nonequilibrium state and allows dynamic control and reconfiguration Banerjee et al. (2020). The theory of elasticity can help identify the associated bending energy $E_{b}$ Landau et al. (2012), see Box II.
@@
which is known as the Kratky–Porod model Kratky and Porod (1949).
At the molecular level, cytoskeletal filaments are linear polymers: monomers that are bound together as an unbranched chain. Physically, they can be distinguished by their stiffness. Thermal fluctuations induce random deformations of the filaments; thus, the thermal energy $k_{\mathrm{B}}T$ is a natural scale to classify the ease of bending. The ratio $\lambda_{\mathrm{p}}=\frac{\beta_{\mathrm{f}}}{k_{\mathrm{B}}T}$ is the persistence length of the filament. By comparing $\lambda_{\mathrm{p}}$ with the filament length we can classify filaments as follows.
@@
If $\lambda_{\mathrm{p}}\ll L$ the filament is considered flexible (the filament will exhibit highly coiled configurations); if $\lambda_{\mathrm{p}}\approx L$ the filament is termed semiflexible, coiling and knotting are not present but thermal fluctuations can induce bending Kroy and Frey (1996); if $\lambda_{\mathrm{p}}\gg L$ the filament is stiff, and it appears as a stiff rod. The persistence length can be measured from the fluctuations of the filament. Under thermal fluctuations, the tangent $\bm{t}$ to different parts of the filament performs a random walk. The correlation between two different points will then exhibit an exponential decay $\langle\bm{t}(s+\Delta s)\bm{t}(s)\rangle=e^{-\Delta s/\lambda_{\mathrm{p}}}$.
@@
Keratin $6\times 10^{3}-10^{8}$ Fudge et al. (2003); McKittrick et al. (2012)
@@
Cellulose $138\times 10^{6}$ Nishino et al. (1995)
Bacterial flagella $(0.1-380)\times 10^{6}$ Shen et al. (2022) $(0.1-100)\times 10^{-23}$ Shen et al. (2022)
Motile cilia $(0.01-120)\times 10^{6}$(†) $(0.3-6)\times 10^{-21}$ Okuno and Hiramoto (1979); Ishijima and Hiramoto (1994)
Filamentous cyanobacteria $(0.4-11.2)\times 10^{3}$ Faluweki and Goehring (2022); Boal and Ng (2010) $(16-46)\times 10^{3}$ Faluweki and Goehring (2022) $(1-20)^{-17}$ Faluweki and Goehring (2022); Kurjahn et al. (2024a)
C. elegans (adult) $(80-140)$(*)Backholm et al. (2013) $\sim 10^{-13}$ Backholm et al. (2013)
Eels and snakes $100-1000$ Guo and Mahadevan (2008); Zhang et al. (2021) $10^{-2}-10^{-3}$ Long Jr (1998)
@@
The cytoskeleton is highly dynamic, because its constituents can be assembled and disassembled as the need arises due to permanent energy dissipation. Cytoskeletal filaments are grouped into three main classes: microtubules, actin filaments, and intermediate filaments Fletcher and Mullins (2010), see Fig. 2. The main differences among these three types of filaments are their mechanical stiffness, polarity, and the molecular motors associated with them Fletcher and Mullins (2010). We now discuss these three classes of biofilaments in turn.
@@
$\bullet$ Microtubules are hollow and polar tubular structures, which grow by the addition of tubulin dimers ($8$ nm in length) bound to GTP (guanosine-5’-triphosphate) molecules Brouhard and Rice (2018); Zhou et al. (2023). Microtubules are the stiffest cytoskeletal filaments, see Table 1. They are formed by $\alpha$- and $\beta$-tubulin protein dimers. The protein dimers polymerize into linear chains, called protofilaments, which arranged as a hollow cylinder constitute a microtubule; most eukaryotic cells contain microtubules with 13 protofilaments, characterised by an internal diameter of about $15$ nm and an external diameter of about $25$ nm Boal (2012); Dustin (2012). Microtubules perform crucial functions in the life of a cell. During cell division, microtubule nucleate from the centrioles that organize the mitotic spindle Azimzadeh and Marshall (2010). Microtubules are also the building blocks of eukaryotic cilia and flagella Grimstone and Klug (1966); Amos and Klug (1974).
@@
The considerable stiffness of microtubules is reflected in a persistent length $\lambda_{\mathrm{p}}$ of about 5 mm, which can span the entire length of an animal cell Fletcher and Mullins (2010), and a Young’s modulus ranging from one to 50 GPa, depending on their length, or on the formation of network structures Gittes et al. (1993); Sharma and Vershinin (2020), see Table 1. Classic experiments employed thermal fluctuations to measure the flexural rigidity of microtubules $\beta_{\mathrm{f}}=2.2\times 10^{-23}$ N m2 Gittes et al. (1993).
@@
Microtubules are highly dynamic systems, where the cell-regulated nonequilibrium processeses of polymerisation and depolymerisation Dogterom and Leibler (1993); Fygenson et al. (1994) of a microtubule coexist (known as ‘dynamic instability’ Mitchison and Kirschner (1984)) and extend to a large fraction of the microtubules in a cell. The tau protein (short for ‘tubulin associated unit’) maintain the stability of microtubules in the axons of neurons Weingarten et al. (1975), however, when hyperphosphorylation occurs, the microtubule disintegrate and the tau proteins form aggregates (neurofibrillary tangles, which are tangles of tau filaments) that are associated with neurodegenerative diseases Palmqvist et al. (2020); Chen et al. (2024).
@@
A single microtubule never reaches a steady state length but persists in prolonged states of polymerisation and depolymerisation that interconvert infrequently Desai and Mitchison (1997). The cell uses the free energy of GTP hydrolysis to drive the nonequilibrium polymerisation; this free energy is $\approx 7.5$ kcal mol-1 under standard conditions, or $\approx 12.5$ kcal mol-1 in vivo Lehninger et al. (1993); Desai and Mitchison (1997).
@@
Recent measurements have shown that as microtubules grow or shrink, they undergo conformational changes. The GTP-tubulin dimers are initially curved at an angle of $\approx 12^{\circ}$; after binding they straighten but exert a stress on the rest of the filament structure; finally, after GTP hydrolysis the $\alpha$-tubulin shortens by $3$ and releases the stress, which can be harnessed to perform work Coue et al. (1991); Koshland et al. (1988); Brouhard and Rice (2018). These conformational changes in the dimers generate long-range interactions along the microtubules that modulate their growth Brouhard and Rice (2018).

### v2

@@ Abstract
Nature uses elongated shapes and filaments to build stable structures, generate motion, and allow complex geometric interactions. In this Review, we examine the role of biological filaments across different length scales. From the molecular scale, where cytoskeletal filaments provides a robust but dynamic cellular scaffolding, over the scale of cellular appendages like cilia and flagella, to the scale of filamentous microorganisms like cyanobacteria, among the most successful genera on Earth, and even to the scale of elongated animals like worms and snakes, whose motility modes inspire robotic analogues. We highlight the general mechanisms that couple form and function. We discuss physical principles and models, such as classical elasticity and the non-reciprocity of active matter, that can be used to trace unifying themes linking these systems across about nine orders of magnitude in length.
@@ I Introduction
The profound connection between form and function in biology has been quantified at least since D’Arcy Wentworth Thompson’s influential work [1]. Because geometrical features influence and direct biological activity through physical forces, it seems worthwhile to explore the link between geometry and biological function. In this Review, we examine a simple geometrical shape –the filament– which is pervasive in the architecture of living organisms, and discuss the function of filamentous shapes and the physical advantages they offer.
@@
Filamentous shapes permeate every length scale in the biological world. Natural evolution has found that the geometrical properties of long aspect-ratio shapes can be used in a multiplicity of useful ways. The most important filamentous shape is the DNA double helix molecule [2], that carries a self-description of the organism and provides a biomolecular realization of a Turing-von Neumann tape [3]. The biopolymers forming the cytoskeleton provide structural integrity to the cell and form an intracellular transportation system. Cytoskeletal networks generate cytokinesis [4], migration [5, 6], and apical contraction [7], and even generate cytoplasmic streaming in the Drosophila oocyte [8].
@@
At length scales of a few microns, cilia and flagella can generate fluid flow and active motion due to their hydrodynamical interactions [9, 10, 11] and can also synchronize to generate metachronal waves [12, 10]. Microorganisms such as filamentous fungi form complex networks of hyphae following available water and resources, and can move cytoplasm along the network [13]. Filamentous fungi are also the group of fungi most often used in biotechnological processes and include perhaps the most important fungus in the history of humanity: Penicillium rubens, which was used by Alexander Fleming to extract penicillin [14, 15, 16]. The slime mold Physarum polycephalum performs cytoplasmic flow across its network akin to a generalised form of peristalsis, which is optimised when the wavelength of the peristaltic wave is of the order of the size of the network [17]; this filamentous network can even encode memory about nutrient location in its morphology [18]. In the prokaryotic domain, cyanobacteria were among the first microorganisms to straddle the transition from unicellular to multicellular organisms [19, 20], and include numerous filamentous species.
@@
Figure 1: Filamentous shapes in biology. (a) Growing complexity of filaments across orders of magnitude in length, from nucleic acids that store information, to biopolymers that provide structural integrity to cells and form networks spanning the entire cytoskeleton. Flagella and cilia self-assemble from polymers to generate motion. Bacterial cells such as cyanobacteria self-organize into filamentous shapes. Animals, from worms and nematodes to snakes and eels utilize elongated shapes to achieve locomotion and collective behaviour. (b) Conceptual classification of biological function emerging from the three main physical properties of biological filaments: (i) the ratio of persistence length to contour length, which determines whether a filament is flexible, semi-flexible, or rigid; (ii) structural complexity, such as twisting or cross-linking into networks, assembly into cilia carpets, organisation of muscle fibres or even an entire organism; (iii) activity, that is, the nonequilibrium conversion of chemical energy (ATP) into persistent motion. An archetypical example is the motion of kinesin on microtubules.
@@
At much larger length scales, organisms such as worms, snakes, and eels have elongated bodies that grant them ecological advantages. Worms, for example, can form entangled ‘blobs’ that exhibit shear-thinning rheology [21] and reversible self-assembly [22]. The physical properties of single filaments can be combined to create emergent complexity and functionality, such as in networks of cross-linking cytoskeletal filaments [23] or in slime molds that can form more optimal foraging networks than Steiner’s minimal tree [24].
@@
The systems described above span about nine orders of magnitude in size: from nucleic acids in the nanometre range in diameter, to snakes and eels in the metre range. Figure 1(a) shows a schematic representation of different biological filaments across these different length scales, and also illustrates the growth in complexity of biofilaments. Despite this apparent diversity, in this work, we posit that it is possible to identify and discuss general physical features of elongated shapes (from macromolecules to whole organisms) that depend on the geometrical form of filaments and, in turn, influence their biological function.
@@
In this context, we ask the following questions. Can we find similarities in the function of the systems depicted in Fig. 1(a)? For example, how do cytoskeletal filaments generate structural integrity in the cell? How do groups of cilia generate fluid flow, or how do large elongated organisms generate undulatory locomotion? Understanding how physical systems interact via simple rules to generate interesting collective behaviour can and will have novel applications. The condensed-matter motto of ‘More is Different’ proposed by Philip Anderson [25] will be a guiding principle in the present work. For example, a self-organizing robot exhibits behaviour intermediate between solid and liquid with adaptive capabilities such that it can flow over obstacles [26]. This combination of flexibility and structural integrity is reminiscent of a cell’s cytoskeleton, or even of snake locomotion. Each of these filamentous systems can serve as inspiration for new biosynthetic materials [27]. Examples include strain-stiffening metamaterials using PDMS slats that mimic cellular behaviour [28], or the use of carbon nanotubes to study the effect of crowding and filament stiffness on their microscopic dynamics [29, 30].
Can we find some general motifs to unify systems as disparate as those highlighted above? In Fig. 1(b), we propose a conceptual diagram containing three main principles that govern the biological function of filaments based on their physical properties. We identify three main axes: (i) activity: every living organism uses adenosine triphosphate (ATP) as the currency of energy. Active matter is defined as any system that converts energy into persistent or systematic motion [31, 32]. Thus, active matter systems are intrinsically out of equilibrium, and have access to a variety of steady states not allowed in equilibrium systems.
(ii) The ratio of persistence length to contour length is a fundamental physical parameter that determines the type of biological activity of a filament: from the rigid microtubules that provide the scaffolding of the cell to more flexible filaments like actin (or arguably the body of snakes and eels) which are more prone to reconfigurations. Physically, the rigidity and length of different biofilaments determine the class of models used to describe their mechanical and dynamical properties. If we consider the angle $\theta$ between the local tangent and a reference axis, thermal and biological fluctuations will bend a filament randomly, such that $\langle(\Delta\theta)^{2}\rangle=\Delta s/\lambda_{\mathrm{p}}$ where $s$ is the arc length along the filament, and $\lambda_{\mathrm{p}}$ is called the filament’s persistence length, which is determined by the relative strength of the random fluctuations and the filament’s bending rigidity [33] (see Sec. II). On the one hand, if the observation length scale (or total length) is much smaller than $\lambda_{\mathrm{p}}$ a filament can be approximated as a smooth line enabling an elastodynamic approach. In this regime, we can mathematically describe filaments as a curve in three-dimensional (3D) space, whose local orientation changes smoothly along their length. Box I describes the Frenet–Serret equations used in differential geometry to characterize a curve.
@@
On the other hand, at small length scales, such as inside the cell, the polymeric nature of biofilaments is consequential and this character is amplified by the nonequilibrium nature of the fluctuations. For example, variations in polymer stiffness and connectivity of actomyosin networks by actin filament bundling and cross-linking proteins can systematically tune the mechanical response of actin networks from extensile to contractile [34, 35].
[…]
(iii) The complexity of the structural organisation. Filaments can link and bundle together in different ways and achieve novel mechanical functionalities. Single globular actin proteins (G-actins) assemble in filaments (F-actin), which together with cross-linking proteins and a myriad of other molecules form cytoskeletal networks. The fascin bundling protein can form filopodia – tightly packed, parallel filaments of actin that exert protrusive forces in migrating cells [45], whereas the disordered actomysoin cortex can sustain tensive stresses. At larger scales, arrays of cilia can exhibit metachronal waves ( travelling waves of synchronized cilia phase [46, 47] which generate self-propulsion in microorganisms or mucociliary clearance in ciliated epithelia [48]. Multiple cells, in turn, can assemble into simple filaments, such as the Oscillatoria lutea cyanobacteria, or form even more complex filamentous organisms, such as nematodes, eels and snakes. We discuss [Fig. 1(b)] in more detail below.
@@
This Review is organised as follows. Section II discusses the role of biofilaments inside the cell. Section  III reviews the role of filamentous cellular appendages, such as the locomotor function of cilia and flagella. In Sec. IV we discuss the biological properties of filamentous bacteria. Section V deals with macroscopic elongated organisms, such as snakes and eels. Finally, in Sec. VI we outline a roadmap of future work and outstanding questions at this exciting frontier shared by biology and physics, to establish a common framework describing the physics of biological filaments (Fig. 1).
@@ II Biopolymers & cytoskeletal filaments
The cytoskeleton is a hierarchically built, 3D polymer-based scaffold that spans the entire interior of a cell and defines cell morphology [49]. The basic building blocks are few key proteins that assemble together into a broad range of filaments [50]. These filaments together with a host of proteins that bind to the filaments’ sides or ends form a network that provides mechanical stability to the cell [51]. Using only a few building blocks allows great flexibility and economy of materials. For example, cytoskeletal polymers achieve the same rigidity (elastic modulus) of gel-like substances with two orders of magnitude less material [52, 51].
At the molecular level, cytoskeletal filaments are linear polymers: monomers that are bound together as an unbranched chain. Physically, they can be distinguished by their stiffness. Thermal fluctuations induce random deformations of the filaments; thus, the thermal energy $k_{\mathrm{B}}T$, where $k_{\mathrm{B}}$ is Boltzmann’s constant, and $T$ the temperature, is a natural scale to classify the ease of bending. The ratio of bending stiffness $\beta_{\mathrm{f}}$ (see Box II) to thermal energy $\lambda_{\mathrm{p}}=\frac{\beta_{\mathrm{f}}}{k_{\mathrm{B}}T}$, called persistence length, has dimensions of a length, and represents the typical length scale over which fluctuations in the filament’s tangent decay. The persistence length can in fact be measured from the equilibrium fluctuations of the filament. Under thermal fluctuations, the tangent $\bm{t}$ to different parts of the filament performs a random walk, and obeys the diffusion equation. The correlation between two different points will then exhibit an exponential decay $\langle\bm{t}(s+\Delta s)\cdot\bm{t}(s)\rangle=e^{-\Delta s/\lambda_{\mathrm{p}}}$.
By comparing $\lambda_{\mathrm{p}}$ with the filament length $L$ we can classify filaments as follows. If $\lambda_{\mathrm{p}}\ll L$ the filament is considered flexible (the filament will exhibit highly coiled configurations); if $\lambda_{\mathrm{p}}\approx L$ the filament is termed semiflexible, coiling and knotting are not present but thermal fluctuations can induce bending [53]; if $\lambda_{\mathrm{p}}\gg L$ the filament is stiff, and it appears as a stiff rod.
Biological filaments are subject to different types of deformations: extensional and compressional forces and moments that cause the filaments to bend and twist. Any deformation mode costs energy, which is continuously used in the cytoskeleton to maintain the cell in a nonequilibrium state and allows dynamic control and reconfiguration [35]. The theory of elasticity can help identify the associated bending energy $E_{b}$ [54], see Box II.
[…]
@@
which is known as the worm-like chain model or Kratky–Porod model [59].
We can modify the Euler–Bernoulli Eq. (1) (see Box II) and apply it to microscopic filaments that experience viscous drag and whose inertia is negligible [60]. Considering model A dissipative dynamics [61, 62]
$\gamma\frac{\partial w}{\partial t}=-\frac{\delta\mathcal{H}_{b}}{\delta w}+\xi(z,t)\,,$ (3)
@@
$\gamma\frac{\partial w}{\partial t}=-\beta_{\mathrm{f}}\frac{\partial^{4}w}{\partial z^{4}}+\xi(z,t)\,.$ (4)
If the filament is under tension of magnitude $f$, the above equation is modified by the inclusion of a second derivative [66, 67, 68]
$\gamma\frac{\partial w}{\partial t}=-\beta_{\mathrm{f}}\frac{\partial^{4}w}{\partial z^{4}}+f\frac{\partial^{2}w}{\partial z^{2}}+\xi(z,t)\,.$ (5)
@@
Keratin $(0.3-0.4)$ [76] $6\times 10^{3}-10^{8}$ [77, 78]
@@
Cellulose $138\times 10^{6}$ [80]
Bacterial flagella $1-40$ [81] $(0.1-380)\times 10^{6}$ [82] $(0.1-100)\times 10^{-23}$ [82]
Bacterial flag. hook $(10-130)\times 10^{-3}$ [83, 84] $(0.01-1)\times 10^{6}$ [84] $(0.001-300)\times 10^{-26}$ (*) [83, 84]
Motile cilia $(0.7-15)\times 10^{5}$(†) $(0.01-120)\times 10^{6}$(‡) $(0.3-6)\times 10^{-21}$ [85, 86]
Fil. cyanobacteria $(0.4-11.2)\times 10^{3}$ [87, 88] $(16-46)\times 10^{3}$ [87] $(1-20)^{-17}$ [87, 89]
C. elegans (adult) $(80-140)$(**)[90] $\sim 10^{-13}$ [90]
Eels and snakes $100-1000$ [91, 92] $10^{-2}-10^{-3}$ [93]
@@
In addition to elasticity, a fundamental aspect of biological filaments is the fact that they are active, that is, they consume ATP to produce motion. Specifically, the cytoskeleton is highly dynamic: its constituents can be assembled and disassembled as the need arises due to permanent ATP consumption and energy dissipation. The overdamped dynamics of a particle can be described by a Langevin equation
$\dot{\bm{x}}=\bm{F}(\bm{x})+\bm{\xi}(t)\,,$ (6)
where $\bm{F}$ represents the deterministic drift driving the system, and $\bm{\xi}$ is a vector whose components are zero-mean white Gaussian noise with variance $\langle\xi_{i}(t)\xi_{j}(t^{\prime})\rangle=2D_{ij}\delta(t-t^{\prime})$, and $D_{ij}$ is the diffusion tensor. The drift may contain a conservative term in addition to a nonconservative one representing the self-propulsion $\bm{F}(\bm{x})=(\bm{v}-\nabla U)$. By coarse-graining, one can obtain the Fokker–Planck equation for the probability density $P(\bm{x},t)$ written in conservative form
$\frac{\partial P(\bm{x},t)}{\partial t}+\nabla\cdot\bm{J}=0\,,$ (7)
where $\bm{J}=\bm{F}P-\nabla\cdot(\bm{D}P)$ is the probability flux. If the process has reached steady state (i.e., after some time $\frac{\partial P_{\mathrm{ss}}}{\partial t}=0$), by Helmholtz–Hodge decomposition, the drift field can be expressed as the sum of reversible and irreversible parts $\bm{F}=\bm{F}_{\mathrm{rev}}+\bm{F}_{\mathrm{irr}}$, where $\bm{F}_{\mathrm{rev}}=\bm{D}\cdot\nabla\ln P_{\mathrm{ss}}+\nabla\cdot\bm{D}$ and $\bm{F}_{\mathrm{irr}}=\bm{J}_{\mathrm{ss}}/P_{\mathrm{ss}}$. For constant noise, the reversible part can be written as the gradient of a potential (thus is curl-free), while $\nabla\cdot\bm{F}_{\mathrm{irr}}=0$ so it is the curl of some vector [94, 95]. The nonequilibrium probability flux, and in particular the rotational drift $\bm{F}_{\mathrm{irr}}$ is a characteristic and strong measure of how far active matter systems are from equilibrium [96, 97, 98].
@@
Cytoskeletal filaments are grouped into three main classes: microtubules, actin filaments, and intermediate filaments [50], see Fig. 2. The main differences among these three types of filaments are their mechanical stiffness, polarity, and the molecular motors associated with them [50]. We now discuss these three classes of biofilaments in turn.
@@
$\bullet$ Microtubules are hollow and polar tubular structures, which grow by the addition of tubulin dimers ($8$ nm in length) bound to GTP (guanosine-5’-triphosphate) molecules [99, 100]. Microtubules are the stiffest cytoskeletal filaments, see Table 1. They are formed by $\alpha$- and $\beta$-tubulin protein dimers. The protein dimers polymerize into linear chains, called protofilaments, which arranged as a hollow cylinder constitute a microtubule. Most eukaryotic cells contain microtubules with 13 protofilaments, characterised by an internal diameter of about $15$ nm and an external diameter of about $25$ nm [70, 101]. Microtubules perform crucial functions in the life of a cell. During cell division, microtubules nucleate from the centrioles that organize the mitotic spindle [102]. Microtubules are also the building blocks of eukaryotic cilia and flagella [103, 104], see Sec. III.
@@
The considerable stiffness of microtubules is reflected in a persistent length $\lambda_{\mathrm{p}}$ of about 5 mm, which can span the entire length of an animal cell [50], and a Young’s modulus ranging from one to 50 GPa, depending on their length, or on the formation of network structures [72, 79], see Table 1. Classic experiments employed thermal fluctuations to measure the flexural rigidity of microtubules $\beta_{\mathrm{f}}=2.2\times 10^{-23}$ N m2 [72]. The contour length of microtubules is typically about $10~{}\mu$m (but in axons they can even reach $100~{}\mu$m) [105, 106]. Thus, the ratio $\lambda_{\mathrm{p}}/L\gg 1$, indicating a strong structural role. In fact, microtubules are involved in key (re)structural processes of the cell; for example, they provide the tracks for kinesin and dynein-driven transport of organelles and vesicles within the cell [107, 108], and are central components of cell mechanics during tissue morphogenesis [109, 110].
@@
Microtubules are highly dynamic systems, where the cell-regulated nonequilibrium processes of polymerisation and depolymerisation [111, 112] of a microtubule coexist (known as ‘dynamic instability[113]) and extend to a large fraction of the microtubules in a cell. The tau protein (short for ‘tubulin associated unit’) maintain the stability of microtubules in the axons of neurons [114], however, when hyperphosphorylation occurs, the microtubules disintegrate and the tau proteins form aggregates (neurofibrillary tangles, which are tangles of tau filaments) that are associated with neurodegenerative diseases [115, 116].
@@
A single microtubule never reaches a steady state length but persists in prolonged states of polymerisation and depolymerisation that interconvert infrequently [117]. The cell uses the free energy of GTP hydrolysis to drive the nonequilibrium polymerisation; this free energy is $\approx 7.5$ kcal mol-1 under standard conditions, or $\approx 12.5$ kcal mol-1 in vivo [118, 117].
@@
Recent measurements have shown that as microtubules grow or shrink, they undergo conformational changes. The GTP-tubulin dimers are initially curved at an angle of $\approx 12^{\circ}$; after binding they straighten but exert a stress on the rest of the filament structure; finally, after GTP hydrolysis the $\alpha$-tubulin shortens by $3$ and releases the stress, which can be harnessed to perform work [119, 120, 99]. These conformational changes in the dimers generate long-range interactions along the microtubules that modulate their growth [99].